I only have a few hours to answer and send. I would appreciate if you help. Thank you so much.a 1. Let V = {[ :a+c=b+ =b+ d} une and T:V + R with T ([: 1) = a +c. Cd a) Find a basis for the kernel of T. dim(Ker(T)) =? (
Answer #1

T: VR atc dim v= 4- No. of L.I restriction 4-1 - 3 dimv=3 kert {C}ev / [24] = 0? - atc= {[EN] í c=-a & ate=bra a b cd bt d=0dim (IMT) = 3-2 =) dim Imt =) و Fer isomorphism dim uz dim R 3 + 1 So, I us not au Isomarplusm T.